40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  PSL  Not logged in ELOG logo
Entry  Fri Aug 30 18:40:01 2019, anchal, Notes, BEAT, Quick note: BN detector changed to NF1811 
    Reply  Wed Sep 4 14:41:33 2019, anchal, Notes, BEAT, Quick note: BN detector changed to SN101 and other changes 
       Reply  Thu Sep 5 19:52:06 2019, anchal, Notes, BEAT, Quick note: FSS gain reduced, Moku input range changed 
          Reply  Tue Sep 10 19:43:02 2019, anchal, Notes, BEAT, Quick note: Detector changed to NF1811 
             Reply  Mon Sep 16 18:03:33 2019, anchal, Notes, BEAT, Quick note: Detector changed to Sn101 
                Reply  Tue Oct 1 17:28:25 2019, anchal, Notes, BEAT, Quick note: Moku connected with USB 
                   Reply  Tue Oct 15 13:23:16 2019, anchal, Notes, Safety, Quick note: Shutting Down Lasers as Safety Sign is not working 
                      Reply  Tue Oct 15 18:53:00 2019, anchal, Notes, Safety, Quick note: Lasers started again 
                         Reply  Wed Oct 16 19:11:11 2019, anchal, Notes, BEAT, Quick note: Cavities locked; Changed BN Freq Counter Poll Rate 
                            Reply  Fri Oct 18 17:47:49 2019, anchal, Notes, BEAT, Quick note: BN Noise bakc to where it was. 
                               Reply  Mon Nov 4 20:13:44 2019, anchal, Notes, BEAT, Quick note: FSS Loop Gain Changes 
                                  Reply  Wed Jan 22 15:20:00 2020, anchal, Notes, BEAT, Notch improved on FSS RFPDs. Gain values increased as well. 
                                     Reply  Thu Jan 30 10:55:22 2020, anchal, Notes, BEAT, Lowest ever beatnote spectrum today. CTN_Daily_BN_Spec.pdf
                                        Reply  Wed Feb 12 10:01:04 2020, anchal, Notes, BEAT, Increased sampling rate to 125kSa/s; lowest noise in higher frequencies 
Message ID: 2412     Entry time: Wed Sep 4 14:41:33 2019     In reply to: 2408     Reply to this: 2418
Author: anchal 
Type: Notes 
Category: BEAT 
Subject: Quick note: BN detector changed to SN101 and other changes 
  • Beatnote detector changed from NF1811 -> SN101
  • Tuned slightly modematching in North path. Increased by 3 %.
  • NFSS Fastgain increased to 14.
  • Switched off all auxilary equipments: Marconi 2023A, SR780, SR785 and AG4395A.
  • Moku is now directly connected to beatnote detector. (Earlier it was connected through a coupler)
  • Moku firmware update to 1.9 is done. Pymoku is upgraded to 2.8.0.
  • Flag '-a' added in dailyBNspec.py code to switch on 10x attenuation for large signals. This makes input range 10Vpp.
  • The new updates have caused some errors to creep in in remote control of Moku. So daily BN measurement is halted.
ELOG V3.1.3-